### v1

# Physical Effects of Gravitational Waves at Second Order

Abstract
There is currently no rigorous definition of gravitational wave strain at second order in cosmological perturbation theory. The usual association of gravitational waves with transverse and traceless fluctuations of the metric on spatial hypersurfaces becomes ambiguous at second order, as it inherently depends on the spacetime slicing. While this poses no practical issues in linearized gravity, it presents a fundamental problem for secondary gravitational waves, especially notorious for gravitational waves induced by primordial fluctuations. We compute, for the first time, the physical effects of gravitational waves at second order, as measured by geodesic observers that emit and receive electromagnetic signals, thereby settling the debate on gauge ambiguities. We find that the measured gravitational wave strain coincides with the transverse-traceless components in the Newton gauge.
Introduction.— Interest in Gravitational Waves (GWs) has grown steadily after the first detection by the LIGO and Virgo collaborations [1]. Recent hints of a GW background reported by Pulsar Timing Array (PTA) collaborations (NANOGrav, EPTA, PPTA, and CPTA) [2, 3, 4, 5, 6, 7, 8, 9, 10] further stimulated studies on the cosmic production of GWs, see, e.g., Refs. [11, 12, 13, 14, 15, 16]. Among possible sources, GWs induced by primordial fluctuations in the early universe [17, 18, 19, 20, 21, 22, 23, 24], so-called induced GWs, have attracted remarkable attention; see Refs. [25, 26] for recent reviews.
Primordial fluctuations are the seeds of Cosmic Microwave Background (CMB) anisotropies and large-scale structures in the Universe [27]. The same fluctuations induce GWs at second order, which predict a scale-invariant spectrum $\Omega_{\mathrm{GW}}\sim 10^{-24}$ on large scales [24]. It could become significant on smaller scales, usually related to the primordial black hole scenario [28, 29, 30, 31, 32, 33, 34, 35, 15, 36, 37] and serve as a probe of the Universe’s unexplored expansion history [38, 39, 40, 13, 41]. Yet, there are lingering gauge ambiguities in theoretical predictions for induced GWs even today [20, 42].
In linearized gravity, GWs are associated with Transverse and Traceless (TT) fluctuations of the spatial metric [43], commonly called tensor fluctuations. While tensor fluctuations are gauge invariant at linear order, they stop being so at second order in cosmological perturbation theory [20, 21, 22], leading to notorious gauge issues in induced GWs [42, 44, 45]. Solving this matter is of vital importance for reliable theoretical predictions and is also of general fundamental interest in GW physics.
Previous studies found that the predicted induced GW spectrum is identical in the Newton, synchronous, and flat gauges but not in, e.g., the comoving gauge [46, 47, 48, 49, 50, 51, 52] (although dissipative effects alleviate the issue [53]). Ref. [50] used the Newton gauge as a benchmark to classify an approximate gauge independence. Such a choice is based on the expectation that the Newton gauge, which recovers Newtonian physics deep inside the Hubble radius, is suitable to describe observations (see also Refs. [21, 22]). Here we provide proof.
Existing proposals to remedy the gauge issue are based on extrapolations of results from flat spacetime. For instance, Ref. [46] argues that the synchronous gauge is the closest to the TT gauge. Ref. [54] proposes to drop terms that do not behave as expected from propagating radiation based on the Sommerfeld boundary condition. Although these ideas provide some insight, it is unclear whether they also apply to cosmological spacetimes without explicit calculation of the measured strain.
In this Letter, we compute, for the first time, the GW strain measured by a detector at second order. For simplicity, we focus on geodesic clocks [55], where two geodesic observers exchange ElectroMagnetic (EM) signals. Time delays are observables and, therefore, free from gauge ambiguities. We find that the physical effect due to GWs coincides with the TT components in the Newton gauge, thereby settling the gauge issue of induced GWs.
GW measurements and cosmic geodesic clock.— Current GW observations rely on EM signal exchanges [56]. For example, in GW interferometers, a laser beam is emitted, split, bounced off mirrors, and received back. PTAs carefully monitor periodic radio signal emitted by milisecond pulsars in the galaxy. Similarly, one may use spacecraft Doppler tracking [57, 58, 59]. In all these measurements, passing GWs cause a delay in the proper time the detector takes to receive the signal, as well as shift in the signal’s frequency.
The elementary process in such measurements is the emission and reception of photons between two separate time-like observers. Let us call the spacetime trajectories of emitter and receptor respectively as $x_{\rm e}^{\mu}(\tau_{\rm e})$ and $x_{\rm r}^{\mu}(\tau_{\rm r})$, where $\tau$ is the corresponding proper time, and the photon geodesic connecting them as $x_{\gamma}^{\mu}(w)$, where $w$ is the affine parameter. We provide an illustration in Fig. 1.
A photon emitted at initial time $\tau_{\rm ei}$ arrives at the receptor at final time $\tau_{\rm rf}$. For initially synchronized observers, $\tau_{\rm rf}$ is the time delay of the signal. But, irrespective of synchronization, pulses periodically emitted every infinitessimal time step $\Delta\tau_{\rm e}$ are received every $\Delta\tau_{\rm r}$ with redshift $\Delta\tau_{\rm r}/\Delta\tau_{\rm e}\to d\tau_{\rm rf}(\tau_{\rm ei})/d\tau_{\rm ei}=1+z$ [60]. Our goal is to compute variations of the time delay and redshift due to passing GWs at second order in cosmological perturbation theory. Here we show our main results, and leave the detailed derivation to a companion paper [61].
To build our geodesic clock, we solve the time-like geodesics of the observers, and the null geodesic of the photons in a perturbed flat Friedmann–Lemaître–Robertson–Walker (FLRW) Universe, the latter with initial and final boundary conditions at emission and reception, i.e. $x_{\gamma}^{\mu}(w_{\rm i})=x_{\rm e}^{\mu}(\tau_{\rm ei})$ and $x_{\gamma}^{\mu}(w_{\rm f})=x_{\rm r}^{\mu}(\tau_{\rm rf})$. Let us sketch our methodology below.
We solve the system perturbatively in terms of proper reception time, which is gauge invariant. Expanding the boundary condition at reception around a background proper reception time $\tau^{(0)}_{\rm rf}$ allows us to isolate perturbative corrections. For instance, at first order we have that
$x^{(1)\mu}_{{\rm r}}(\tau^{(0)}_{\rm rf})+p^{(0)\mu}_{{\rm r}}(\tau^{(0)}_{\rm rf})\,\tau^{(1)}_{\rm rf}=x^{(1)\mu}_{{\gamma}\rm f}\,$, (1)
where $p^{\mu}_{{\rm r}}(\tau_{\rm r})\equiv dx^{\mu}_{{\rm r}}/d\tau_{\rm r}$ is the receptor’s 4-momentum and, here and henceforth, we use a superscript $(l)$ to denote the order in the perturbation expansion. With the null geodesic solution, the zeroth component of Eq. (1) links $\tau^{(1)}_{\rm rf}$ to $\tau_{\rm ei}$ and metric perturbations in between. We proceed in a similar fashion at second order.
Figure 1: [Lower] Illustration of the elementary process. Time flows upwards. Black lines depict the time slices, blue lines show the world lines of the observers and red lines show the propagating light rays. Dashed and solid lines respectively correspond to the background and perturbed trajectories. [Upper] Zoom-in of reception showing first-order corrections as in Eq. (1).
In conformal coordinates $x^{\mu}=(\eta,x^{i})$, the perturbed FLRW metric is [62]
${\rm d}s^{2}=$ $g_{\mu\nu}dx^{\mu}dx^{\nu}=a^{2}(-e^{2\phi}{\rm d}\eta^{2}$
$\quad\,+e^{2\psi}\big(e^{Y}\big)_{ij}({\rm d}x^{i}+N^{i}{\rm d}\eta)({\rm d}x^{j}+N^{j}{\rm d}\eta)).$ (2)
In the scalar-vector-tensor decomposition [63, 64, 65], and ignoring vector fluctuations for simplicity, we have that $N^{i}=\delta^{ij}\partial_{j}\beta$ and $Y_{ij}=2(\partial_{i}\partial_{j}-\delta_{ij}\partial^{2}/3)E+h_{ij}$, with $h_{ij}$ the TT components usually associated with GWs. Note that we exclusively use $\partial^{2}\equiv\delta^{ij}\partial_{i}\partial_{j}$ for the Laplacian.
A subtle issue is the initial conditions. To keep it general, we only fix the relative spacetime background locations of emitter and receiver at rest. Namely, we assume they are separated in a direction $n^{i}$ by a comoving distance $L$, that is $x^{(0)\mu}_{{\rm r}}-x^{(0)\mu}_{{\rm e}}=(0,L\,n^{i})$. We leave $n^{i}$ free as one could imagine a sphere of detectors exchanging EM signals in all directions. Fully fixing initial conditions requires specifying the synchronization’s time-slice. However, one may consider a situation with no initial spacetime fluctuations around the detector, only to be perturbed after emission. Time delays due to passing GWs must be independent of such initial uncertainty.
First-order results.— Let us take a closer look at first-order corrections, since they also appear at second-order. After integrating the null condition, i.e. $k^{\mu}_{{\rm\gamma}}k_{{\rm\gamma}\mu}=0$ with $k^{\mu}_{{\rm\gamma}}\equiv dx_{\gamma}^{\mu}/dw$ the photon’s 4-momentum, we calculate the first-order time delay in (1) by integrating by parts and using boundary conditions, which reads
${\tau^{(1)}_{\rm rf}}=a_{\rm rf}\Big($ $\int_{0}^{L}\hskip-0.28453pt{\rm d}\lambda\,\Big(\frac{1}{2}\,n^{i}n^{j}h_{ij}^{(1)}-\Phi^{(1)}-\Psi^{(1)}\Big)$ (3a)
$-X^{(1)0}_{\rm rf}+X^{(1)0}_{\rm ei}+n_{i}(X^{(1)i}_{\rm rf}-X^{(1)i}_{\rm ei})\Big)\,$, (3b)
where $\Phi^{(1)}$ and $\Psi^{(1)}$ are the (gauge invariant) Bardeen potentials [65] and we defined ${\rm d}\lambda=dw/a^{2}(w)$. Fixing $\lambda_{\rm i}=0$, the background solution gives $\lambda_{\rm f}=L$. We also introduced the gauge invariant spacetime locations,
$X^{(1)0}=x^{(1)0}-\sigma^{(1)},\quad X^{(1)i}=x^{(1)i}+E^{(1)}_{,i}\,$, (4)
with $\sigma=\beta-E^{\prime}$ the spacetime shear.
The time delay (3) is manifestly gauge invariant. The first line (3a) contains the well-known GW and Shapiro time delay, respectively. The second line (3b) includes the unspecified initial conditions and the observer’s motion effects. This is clearer if we split the receptor’s location into initial condition and geodesic evolution, namely $x^{(1)\mu}_{\rm rf}=x^{(1)\mu}_{\rm ri}+\Delta x^{(1)\mu}_{\rm r}$, where $\Delta x^{(1)\mu}_{\rm r}=\int_{\eta_{\rm i}}^{\eta_{\rm f}}d\eta\,ap^{(1)\mu}_{\rm r}$. Doing so, we recast the second line (3b) as
${\tau^{(1)}_{\rm rf}}\supset a_{\rm rf}$ $\Big(-\Delta T_{\rm i}+\Delta L_{\rm i}+\Big(1-\frac{a_{\rm ri}}{a_{\rm rf}}\Big)X^{(1)0}_{\rm ri}$ (5a)
$\quad+\int_{\eta_{\rm i}}^{\eta_{\rm f}}d\eta\,n^{i}V^{(1)}_{{\rm r},i}+\frac{1}{a_{\rm rf}}\int_{\eta_{\rm i}}^{\eta_{\rm f}}d\eta\,a\Phi^{(1)}\Big)\,$, (5b)
where we used that $v_{{\rm r},i}=p^{(1)i}_{{\rm r}}/p^{(0)}_{{\rm r}0}$ and introduced the (gauge invariant) velocity potential, $V=v+E^{\prime}$. We also defined $\Delta T_{\rm i}=X^{(1)0}_{\rm ri}-X^{(1)0}_{\rm ei}$ and $\Delta L_{\rm i}\equiv n_{i}(X^{(1)i}_{\rm ri}-X^{(1)i}_{\rm ei})$ for later convenience. At the background we have that $\eta_{\rm f}-\eta_{i}=L$. The solution for $V_{{\rm r}}$ can be found in the End Matter. Noting that the proper time is given by $d\tau_{\rm r}=d\eta/p^{0}_{\rm r}$ and so $\tau_{\rm rf}^{(0)}=\int_{\eta_{\rm i}}^{\eta_{\rm f}}a(\eta)d\eta$, we conclude that the terms in (5a) correspond to a redefinition of $\tau_{\rm rf}^{(0)}$ via
$\eta_{\rm i}\to\eta_{\rm i}+X^{(1)0}_{\rm ri}\,\,\,,\,\,\,\eta_{\rm f}\to\eta_{\rm f}+X^{(1)0}_{\rm ri}-\Delta T_{\rm i}+\Delta L_{\rm i}\,.$ (6)
For the receptor, the signal left $\Delta T_{\rm i}$ earlier and traveled $\Delta L_{\rm i}$ more so it arrived $\Delta T_{\rm i}$ earlier and $\Delta L_{\rm i}$ later, in addition to a shift of the receptor’s conformal time by $X^{(1)0}_{\rm ri}$. The remaining terms in (5b) are due to the observer’s motion, respectively due to the velocity and time delay gained from falling in a gravitational potential $\Phi^{(1)}$.
Despite being gauge invariant, the time delay (3) is not very useful without a complete knowledge of the initial state of the geodesic clock. Although out of the scope of this paper, one could imagine a setup where the initial separation and proper motion of emitter and receiver are known by other means. In that case, one could isolate the propagation effects in the time delay.
To remove initial ambiguities, it is more appropriate to consider the variations in the time delay from successive infinitessimally close pulses, namely the redshift. Using that $1+z=d\tau_{\rm rf}(\tau_{\rm ei})/d\tau_{\rm ei}$, we find that the first order correction reads
$\frac{z^{(1)}}{1+\tilde{z}}$= $\,\Phi^{(1)}_{\rm rf}-\Phi^{(1)}_{\rm ei}-\int_{0}^{L}{\rm d}\lambda\,\big({\Phi^{(1)}}^{\prime}+{\Psi^{(1)}}^{\prime}\big)$
$\quad\,+n^{i}(V^{(1)}_{{\rm r},i}-V^{(1)}_{{\rm e},i})+\int_{0}^{L}{\rm d}\lambda\,n^{i}n^{j}{h^{(1)}_{ij}}^{\prime}\,$, (7)
where we introduced the measured redshift $1+\tilde{z}\equiv{a_{\rm r}(\tau_{\rm r})}/{a_{\rm e}(\tau_{\rm e}})$ [66], with $\tau$ being the non-linear proper time of emitter and observer. This results coincide with the conventional definition, that is $1+z=(p_{\mu}k_{\gamma}^{\mu})_{\rm r}/(p_{\mu}k_{\gamma}^{\mu})_{\rm e}$, and contains the standard and integrated Sachs-Wolfe, the Doppler and the GW effects [67].
Second-order results.— From Eqs. (3a) and (Physical Effects of Gravitational Waves at Second Order), we see that the GW effect comes from the term proportional to $n^{i}n^{j}$. As general second-order calculations are quite involved (see, e.g., Refs. [68, 69, 70, 71]), we focus solely on terms containing $n^{i}n^{j}$ from now on. These terms lead to the characteristic “quadrupole” nature of GWs.
We now proceed as in the first-order case. We expand the boundary condition $x_{\gamma}^{\mu}(w_{\rm f})=x_{\rm r}^{\mu}(\tau_{\rm rf})$ up to second order, for coordinates and proper time, and use the integrated null condition with the solutions to the first order null geodesics to relate $\tau^{(2)}_{\rm rf}$ to the initial boundary condition. We provide more details in the End Matter.
After integration by parts, we find that quadrupole component of the second-order time delay is given by
$\tau^{(2)}_{\rm rf}\supset a_{\rm rf}$ $\,n^{i}n^{j}\Bigg(\frac{1}{2}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}+V^{(1)}_{{\rm rf},i}\Delta X^{(1)}_{j}$
$\quad\qquad+\left(\frac{a^{\prime}_{\rm rf}}{2a_{\rm rf}^{2}}-\frac{a_{\rm rf}^{4}}{2w_{\rm f}^{2}}L\right)\Delta X^{(1)}_{i}\Delta X^{(1)}_{j}\Bigg)\,$, (8)
where we defined $\Delta X^{(1)}_{i}=\int_{0}^{L}d\eta\,V^{(1)}_{{\rm r},i}$ and we dropped terms proportional to the initial displacements from (5a). $h^{N(2)}_{ij}$ is the gauge invariant combination which coincides with the TT components in Newton gauge, that is [62]
[…]
Eq. (Physical Effects of Gravitational Waves at Second Order) is the main result of our work. It explicitly shows that the detector response to passing GWs is sensitive to the gauge invariant combination $h^{N(2)}_{ij}$. In other words, theoretical predictions for GWs are more conveniently derived in the Newton gauge, where $h_{ij}$ coincides with $h^{N(2)}_{ij}$, which is what a GW detector measures. Let us emphasize that our formalism is gauge invariant and that we did not fix any gauge. It so happened that propagation effects in the time delay of a geodesic clock are directly proportional to $h^{N(2)}_{ij}$.
The quadrupole component in the time delay (Physical Effects of Gravitational Waves at Second Order) also contain boundary terms proportional to $\Delta X^{(1)}_{j}$. These terms are exclusively due to the observer’s motion. For instance, some of this terms are proportional to $n^{i}n^{j}V_{{\rm r},i}V_{{\rm r},j}$. If we consider our sphere of receptors, such boundary terms originate from dynamical deformations of the initial spherical configuration. But they have nothing to do with passing GWs. Thus, if one keeps track of the receptors motions by independent means, e.g., via additional exchange of signals between receptors, one can isolate the propagation effects and measure $h^{N(2)}_{ij}$. However, as in the first-order case, making use of time delay requires extensive knowledge of the initial configuration.
Our results can be applied to more general situations. Again, it is intriguing to start from the first-order results. Since time delay is subject to initial ambiguities, let us consider the exchange of pulses and the resulting redshift (Physical Effects of Gravitational Waves at Second Order). Each term in (Physical Effects of Gravitational Waves at Second Order) exhibits a distinct pattern, which can be used to disentangle the different effects. For example, in the context of PTAs, we can first average over time and over pulsars at different positions. This average is insensitive to propagation effects. Then, one can reconstruct the gravitational potential to account for the time-dilation effects, that is those associated with the expansion and the acceleration induced by the gravitational potential. This ideal procedure allows us to isolate the propagation effect at different times.
To compare with PTA literature, consider a coherent plane GW along the $x=x^{1}$ direction given by $h^{(1)}_{ij}=h(t-x)e^{+/\times}_{ij}$, where $e^{+/\times}_{ij}$ are the ++ or $\times$ polarization tensors. Using Eq. (Physical Effects of Gravitational Waves at Second Order), we recover the well-known GW induced redshift [57, 72], namely
$z^{(1)}_{\rm GW}=\,\frac{1}{2}\frac{n^{i}n^{j}e^{+/\times}_{ij}}{1-n^{x}}\big(h(t-L\,n^{x})-h(t)\big)\,$, (10)
where for PTAs we approximated ${a_{\rm rf}}\approx{a_{\rm ei}}$. The accumulative effect of the redshift on the pulse arrival time are called the timing residual. Note that the timing residual is equal to the time delay (3) due to GWs up to a time-independent strain. In PTAs, angular correlations of the GW-induced redshift follow the so-called Hellings–Downs angular correlation; a direct consequence of the quadrupolar nature of GWs [73].
Similarly, from the second-order time-delay (Physical Effects of Gravitational Waves at Second Order), we conclude that the redshift induced by passing GWs at second-order is given by
$z^{(2)}_{\rm GW}$ $=\frac{a_{\rm rf}}{a_{\rm ei}}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,{h^{N(2)}_{ij}}^{\prime}$. (11)
This shows that it is the second-order TT component in the Newton gauge which directly maps to the GW effect. Our results demonstrate that the GW strain power spectrum in Newton gauge is the physical, i.e. observable, quantity for induced GWs. With this, we settle the debate on the gauge issue of induced GWs.
At this point, let us discuss implications of our results on the time-independent GW strain induced by the constant gravitational potential $\Phi^{(1)}$ in matter dominated universes [74, 24, 75, 42, 50, 76, 77]. This constant solution, if generated in early matter dominated phases, has been shown to be very sensitive to the transition to the radiation dominated phase [78, 79]. Looking at Eq. (11), it is clear that this constant mode does not affect the redshift. Therefore, GW detectors using time series (like PTAs and Doppler tracking) will not see such a time-independent strain. The same applies to CMB temperature fluctuations and B-modes [80, 81, 82, 83] as they also depend on the redshift. Furthermore, it does not affect the GW energy density if we consider a definition in which $\rho_{\rm GWs}\propto\langle h_{ij}^{\prime}h_{ij}^{\prime}\rangle$ as proposed by Refs. [51, 77].
However, despite not propagating as a GW, one may wonder whether one could measure such a time-independent strain at all. In fact, such constant strain also appears in the Post-Newtonian approximation [21]. Our interpretation, based on our results of the second-order time delay (Physical Effects of Gravitational Waves at Second Order), is that such a constant strain behaves like the GW memory effect [84, 85]. For instance, once the strain becomes constant, it would appear as
$\tau^{(2)}_{\rm rf}\supset a_{\rm f}\,$ $\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}(x^{i})$
$=a_{\rm f}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}n_{k}{\rm d}x^{k}\,h^{N(2)}_{ij}(x^{i})\,$, (12)
where we used that since $h^{N(2)}_{ij}(x^{i})$ does not depend on time $\int d\lambda=\int n_{k}{\rm d}x^{k}$. The resulting effect is that our sphere of detectors is now displaced by $\Delta L_{\rm i}(x^{i})=\frac{1}{2}n^{i}n^{j}h^{N(2)}_{ij}(x^{i})$; see Eq. (6) and the surrounding discussion. If we calibrate the detector after generation of the constant strain, we eliminate the memory effect. However, if our geodesic clock was calibrated before strain generation, we should be able to see the strain build-up as in the GW memory effect. Moreover, since the constant strain produces a non-trivial Riemann tensor, it may have non-local effects. Thus, despite Ref. [50] showed that the constant mode can be locally gauged away, it is conceivable that it could be observed by extended measurements.
Conclusion.—In this work, we computed the time delay caused by passing GWs at second order in cosmological perturbation theory as measured by geodesic observers exchanging EM signals, without gauge fixing. We showed that the time delay measured by a GW detector is gauge invariant. Then, we identified the effect of passing GWs as that with a quadrupolar dependence on the direction of propagation of the EM signal. We found that, after removing ambiguities associated to the motion of the observers, the resulting physical GW effect to the time delay at second order coincides with the TT components of the spatial metric evaluated in the Newtonian gauge. Therefore, the gauge issue of the induced GW strain is solved.
Note that, although we assumed geodesic observers, which is an appropriate approximation for PTAs and space-borne interferometers like LISA, our calculations can be generalized to non-geodesic observers like LIGO, by considering the acceleration caused by external forces. However, external forces only affect the subtraction associated with the observer’s motion but not the effect due to propagating GWs.
We have only focused on scalar squared contributions to the second-order transverse-traceless components, since they are the relevant quantities for the (scalar-scalar) induced GWs. But in general, there are also scalar-tensor contributions. In fact, it has been found that scalar-tensor induced GW suffer from divergences in some cases [86, 87, 88, 89]. If it is a gauge issue, it would be clear in a generalization of the time delay calculation to include scalar-tensor terms. It would also be interesting to study the geodesic deviation equation in terms of local tetrads and confirm that $h_{ij}^{N(2)}$ (9) is the quantity describing the effect of GWs at second order.
We conclude by clarifying that, although we demonstrated that the GW strain measured by a GW detector coincides with one in the Newton gauge, our result does not imply that the energy density associated with GWs at second order coincides with the one computed in the Newton gauge. Namely, there is an additional gauge ambiguity in the definition of the (pseudo) energy momentum tensor of GWs [90] at second order in cosmological perturbation theory (see, e.g., Ref. [51] for a study in a fixed time-slice). This is, however, a different problem. It concerns the backreaction of second-order perturbations on the background spacetime, which is important for estimating the expansion rate of the Universe, but is unrelated to the strain measured by a GW detector. We leave this issue for future work.
Acknowledgements.—We thank Domenico Giulini, Misao Sasaki and Cheng-jun Fang for valuable discussions. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, and by the DFG under the Emmy-Noether program project number 496592360. G.D. and S.P. also acknowledge support by JSPS KAKENHI grant No. JP24K00624. S.P. and A.W. are supported in part by the National Natural Science Foundation of China (NSFC) Grants Nos. 12475066 and 12447101. A.W. is also supported by the UCAS Joint PhD Training Program.

### v2

@@
 The same fluctuations induce GWs at second order, which predict a scale-invariant spectrum $\Omega_{\mathrm{GW}}\sim 10^{-24}$ on large scales [24].
-It could become significant on smaller scales, usually related to the primordial black hole scenario [28, 29, 30, 31, 32, 33, 34, 35, 15, 36, 37] and serve as a probe of the Universe’s unexplored expansion history [38, 39, 40, 13, 41].
+It could become significant on smaller scales, usually related to the primordial black hole scenario [28, 29, 30, 31, 32, 33, 34, 35, 15, 36, 37, 38] and serve as a probe of the Universe’s unexplored expansion history [39, 40, 41, 13, 42].
 Yet, there are lingering gauge ambiguities in theoretical predictions for induced GWs even today [20, 43].
@@
 where we introduced the measured redshift $1+\tilde{z}\equiv{a_{\rm r}(\tau_{\rm r})}/{a_{\rm e}(\tau_{\rm e}})$ [67], with $\tau$ being the non-linear proper time of emitter and observer.
-This results coincide with the conventional definition, that is $1+z=(p_{\mu}k_{\gamma}^{\mu})_{\rm r}/(p_{\mu}k_{\gamma}^{\mu})_{\rm e}$, and contains the standard and integrated Sachs-Wolfe, the Doppler and the GW effects [67].
+This results coincide with the conventional definition, that is $1+z=(p_{\mu}k_{\gamma}^{\mu})_{\rm r}/(p_{\mu}k_{\gamma}^{\mu})_{\rm e}$, and contains the standard and integrated Sachs-Wolfe, the Doppler and the GW effects [68, 69].
 Second-order results.— From Eqs.
@@
 As general second-order calculations are quite involved (see, e.g., Refs.
-[68, 69, 70, 71]), we focus solely on terms containing $n^{i}n^{j}$ from now on.
+[70, 71, 72, 73, 74, 75, 76, 77]), we focus solely on terms containing $n^{i}n^{j}$ from now on.
 These terms lead to the characteristic “quadrupole” nature of GWs.
@@
 Furthermore, it does not affect the GW energy density if we consider a definition in which $\rho_{\rm GWs}\propto\langle h_{ij}^{\prime}h_{ij}^{\prime}\rangle$ as proposed by Refs.
-[51, 77].
+[90, 52, 83].
 However, despite not propagating as a GW, one may wonder whether one could measure such a time-independent strain at all.
@@
 For instance, once the strain becomes constant, it would appear as
-$\tau^{(2)}_{\rm rf}\supset a_{\rm f}\,$ $\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}(x^{i})$
-$=a_{\rm f}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}n_{k}{\rm d}x^{k}\,h^{N(2)}_{ij}(x^{i})\,$, (12)
+$\tau^{(2)}_{\rm rf}\supset a_{\rm rf}\,$ $\frac{1}{2}n^{i}n^{j}\int_{0}^{L}{\rm d}\lambda\,h^{N(2)}_{ij}(x^{i})$
+$=a_{\rm rf}\,\frac{1}{2}n^{i}n^{j}\int_{0}^{L}n_{k}{\rm d}x^{k}\,h^{N(2)}_{ij}(x^{i})\,$, (12)
 where we used that since $h^{N(2)}_{ij}(x^{i})$ does not depend on time $\int d\lambda=\int n_{k}{\rm d}x^{k}$.
-The resulting effect is that our sphere of detectors is now displaced by $\Delta L_{\rm i}(x^{i})=\frac{1}{2}n^{i}n^{j}h^{N(2)}_{ij}(x^{i})$; see Eq.
+The resulting effect is that our sphere of detectors is now displaced by $\Delta L_{\rm i}(x^{i})=\frac{1}{2}\int_{0}^{L}n_{k}{\rm d}x^{k}n^{i}n^{j}h^{N(2)}_{ij}(x^{i})$; see Eq.
 (6) and the surrounding discussion.
@@
 Then, we identified the effect of passing GWs as that with a quadrupolar dependence on the direction of propagation of the EM signal.
-We found that, after removing ambiguities associated to the motion of the observers, the resulting physical GW effect to the time delay at second order coincides with the TT components of the spatial metric evaluated in the Newtonian gauge.
+We found that, after removing ambiguities associated with the motion of the observers, the resulting physical GW effect on the time delay at second order coincides with the TT components of the spatial metric evaluated in the Newton gauge.
 Therefore, the gauge issue of the induced GW strain is solved.
+Our result also applies to a more general setting, as we have not made any assumptions about the GW source.
+Namely, we have demonstrated that GW strain predictions from any cosmic source, performed in the Newton gauge, are free from gauge artifacts and directly related to observables.
 Note that, although we assumed geodesic observers, which is an appropriate approximation for PTAs and space-borne interferometers like LISA, our calculations can be generalized to non-geodesic observers like LIGO, by considering the acceleration caused by external forces.
@@
 But in general, there are also scalar-tensor contributions.
-In fact, it has been found that scalar-tensor induced GW suffer from divergences in some cases [86, 87, 88, 89].
+In fact, it has been found that scalar-tensor induced GW suffer from divergences in some cases [93, 94, 95, 96, 97].
 If it is a gauge issue, it would be clear in a generalization of the time delay calculation to include scalar-tensor terms.
@@
 Namely, there is an additional gauge ambiguity in the definition of the (pseudo) energy momentum tensor of GWs [98] at second order in cosmological perturbation theory (see, e.g., Ref.
-[51] for a study in a fixed time-slice).
+[99, 52] for a study in a fixed time-slice).
 This is, however, a different problem.